OpenStudy (marco26):

6. The tangent and a secant to a circle are from the same external point. If the tangent is 6 inches and the external segment of the secant is 3 inches, compute the length of the secant.

OpenStudy (zpupster):

If a secant segment and tangent segment are drawn to a circle from the same external point, the product of the length of the secant segment and its external part equals the square of the length of the tangent segment. b*c=a^2 plugging in 3*c=6^2

OpenStudy (marco26):

Thank you :)